Torsion analytique é equivariante sur ℙ^{n}ℂ

Kai KÖ OHLER

Rubrique: Géeoméetrie difféerentielle (19)

Abstract: We calculate an equivariant version of the complex Ray-Singer torsion for all bundles on the ℙ^{1}ℂ and for the trivial line bundle on ℙ^{n}ℂ, for isometries which have isolated fixed points. The result gives for all n a part of the Gillet-Soulée R-genus.

Réesumée: On calcule une version éequivariante de la torsion complexe de Ray- Singer pour tous les fibrées sur ℙ^{1}ℂ et pour le fibrée trivial sur ℙ^{n}ℂ relativement à a une isoméetrie à a points fixes isolées. Le réesultat donne pour tous les n une partie du genre R de Gillet-Soulée.

Version franç caise abré egé ee

La torsion analytique a éetée construite par Ray et Singer [11] comme un analogue analytique de la torsion de Reidemeister. Partant d’un réesultat de Quillen, Bismut, Gillet et Soulée [1] ont prouvée de remarquables propriéetées pour la torsion des fibrées vectoriels sur des fibrations. Soit en effet π : M → B une application holomorphe propre de variéetées complexes, et soit ξ un fibrée vectoriel holomorphe hermitien sur M. Soit Rπ_{∗}ξ l’image directe de ξ. Alors la torsion analytique des fibres de π induit une méetrique sur le déeterminant de la cohomologie de Knudsen-Mumford λ^{KM} := (det Rπ_{∗}ξ)^{−1} qui est un fibrée en droites holomorphe sur B. Dans [1], on a calculée la courbure de la connexion holomorphe hermitienne de la méetrique de Quillen. On obtiens ainsi une version raffinéee du théeorèeme de Riemann-Roch-Grothendieck.

Soit maintenant i : Y ↪ X une immersion de variéetées compactes complexes. Soit η un fibrée vectoriel holomorphe hermitien sur Y , et soit ξ une réesolution de η par un complexe des fibrées vectoriels sur X. Bismut et Lebeau [2] ont calculée la relation entre les méetriques de Quillen sur les déeterminants de la cohomologie de η et de ξ. A l’aide de ces réesultats, Gillet et Soulée [6] ont déemontrée un théeorèeme de Riemann-Roch en géeoméetrie d’Arakelov pour la premièere classe de Chern de l’image directe (voir [12]). Ce théeorèeme a éetée éetendu par Faltings [3] à a tout le caractèere de Chern.

La preuve du théeorèeme de Riemann-Roch utilise un calcul par Gillet, Soulée et Zagier [5] de la torsion sur les espaces projectifs complexes. Ce calcul a permis à a Gillet et Soulée de conjecturer un théeorèeme de Riemann-Roch. Ce mêeme calcul assez difficile donne en particulier le genre R de Gillet-Soulée, qui apparaîıt explicitement dans le théeorèeme de Riemann-Roch. C’est le genre additif associée à a la séerie

R(x) = ∑_{\begin{matrix}ℓ≥1 \\ odd \end{matrix}} (2ζ′(−ℓ) + ζ(−ℓ) ∑_{j=1}^{ℓ} \frac{1}{j})\frac{x^{ℓ}}{ℓ!} ,

ou ζ est la fonction zêeta de Riemann.

Considéerons maintenant une isoméetrie holomorphe d’un fibrée vectoriel hermitien sur une varietée kä ahlerienne. On peut déefinir une version éequivariante de la torsion en correspondance avec le calcul de Ray de la torsion analytique réeelle pour des espaces lenticulaires. Soit en effet M une variéetée kä ahlerienne de fibrée tangent holomorphe T M, soit ξ un fibrée vectoriel hermitien sur M et soit □_{q} l’opéerateur de Laplace-Kodaira agissant sur des sections de Λ^{q}T^{∗(0,1)}M ⊗ ξ. Soit g une isoméetrie holomorphe de M. Supposons que le fibrée et les méetriques soient invariants par l’action de g. Soit Eig_{λ}(□_{q}) l’espace propre de □_{q} correspondant à a la valeur propre λ et g^{∗} l’action induite par g sur Γ(Λ^{q}T^{∗(0,1)}M ⊗ ξ). Considéerons la fonction

Z(g, s) := ∑_{\begin{matrix}q>0 \\ λ∈ Spec⊓ ⊔_{q} \\ λ≠0 \end{matrix}}(−1)^{q+1}qλ^{−s} Trg_{| Eig_{λ}(□_{q})}^{∗} ,

pour Re s ≫ 0. La torsion éequivariante de M relative à a l’action de g est alors déefinie comme l’exponentielle de la déerivéee à a zéero Z′(g, 0) du prolongement holomorphe de Z(g,.),

τ(g) := e^{− \frac{1}{2} Z′(g,0)} .

Dans cet article on préesente le calcul de la torsion éequivariante pour les fibrées holomorphes sur ℙ^{1}ℂ et pour le fibrée trivial sur les ℙ^{n}ℂ relativement à a une isoméetrie à a points fixes isolées, pour la méetrique de Fubini-Study. Pour une rotation d’angles ∈ π . ℚ, on obtient une expression ferméee oú u la fonction gamma intervient. Pour des angles arbitraires,le resultat final fait intervenir une partie de la fonction R de Gillet-Soulée [5].

Cette partie s’obtient comme suit. Soit pour 0 < φ < 2π et s > 0 ζ^{rot}(φ, s) la séerie Dirichlet

ζ^{rot}(φ, s) := ∑_{k≥1} \frac{sin kφ}{k^{s}} .

Alors ζ^{rot} peut êetre vu comme la partie imaginaire d’une fonction ζ de Lerch. Soit R^{rot}(φ) := \frac{∂}{∂s}ζ^{rot}(φ, 0). On obtient aiséement en utilisant des réesultats classiques la

Proposition 1. R^{rot} est donnée par la formule

R^{rot}(φ) = \frac{C + ln φ}{φ} − ∑_{{ℓ≥1 \atop ℓ odd}} ζ′(−ℓ)(−1)^{\frac{ℓ+1}{2}} \frac{φ^{ℓ}}{ℓ!} .

Si φ = 2π \frac{p}{q} avec p, q ∈ ℕ, 0 < p < q, alors

R^{rot}(φ) = −\frac{1}{2} ln q. cot \frac{φ}{2} + ∑_{ℓ=1}^{q−1} ln Γ (\frac{j}{q} ). sin jφ .

Soit E := O(k_{1}) ⊕ … ⊕ O(k_{n}) un fibrée vectoriel holomorphe sur ℙ^{1}ℂ, éequipée de la méetrique standard (i.e. la courbure de O(1) est la forme de Fubini-Study). D’aprèes un théeorèeme de Grothendieck [4], tout fibrée vectoriel holomorphe sur ℙ^{1}ℂ possèede cette forme. On trouve le Thé eorè eme 2. La torsion analytique éequivariante τ(E, φ) relativement à a une rotation d’un angle φ ∈]0, 2π[ est donnéee par la formule

−2 ln τ(E, φ) = \frac{2R^{rot}(φ)}{sin \frac{φ}{2}}.∑_{j=1}^{n} cos(k_{j}+1)\frac{φ}{2}+∑_{j=1}^{n} ∑_{m=1}^{|k_{j}+1|} \frac{sin(2m − |k_{j} + 1|)\frac{φ}{2}}{sin \frac{φ}{2}} ln j .

Soit maintenant Φ := ( ..._{0  iφ_{n+1}}^{iφ_{1}  0} ) un éeléement de la sous-algèebre

Cartan maximale (canonique) de 𝔰𝔲(n + 1), déefinissant une isoméetrie infinitéesimale de ℙ^{n}ℂ ≅ SU(n + 1)/S(U(1) × U(n)). Supposons que les φ_{j} sont distincts. Alors on a le

Thé eorè eme 3. La torsion éequivariante τ(O, e^{Φ}) pour le fibrée trivial en droite O sur ℙ^{n}ℂ est donnéee par

−2 ln τ(O, e^{Φ}) = (−1)^{n} ∑_{{j,k=1 \atop j≠k}}^{n+1} 2iR^{rot}(φ_{j}−φ_{k}) ∏_{{ℓ=1 \atop ℓ≠k}}^{n+1}(e^{i(φ_{k}−φ_{ℓ})}−1)^{−1}− ln n! .

English version

The analytic torsion was constructed by Ray and Singer [11] as an analytic analogue to the Reidemeister torsion. Bismut, Gillet and Soulée

[1] proved as an extension of a result of Quillen remarkable properties of the torsion in connection with vector bundles on fibrations:

Let π : M → B be a proper holomorphic map of complex manifolds and let ξ be a hermitian holomorphic vector bundle on M. Let Rπ_{∗}ξ be the right-derived direct image of ξ. Then the analytic torsion of the fibres of π induces a metric on the Knudsen-Mumford determinant λ^{KM} := (det Rπ_{∗}ξ)^{−1} which is a holomorphic line bundle on B. The curvature of this Quillen metric as well as its behaviour under changes of the metrics on M and ξ was expressed in [1] explicitly by means of secondary Bott-Chern classes. In particular this gives a refinement of the Riemann-Roch theorem for families.

On the other hand let i : Y ↪ X be an embedding of compact complex manifolds. Let η be a hermitian holomorphic vector bundle on Y and let ξ be a resolution of η by a complex of vector bundles on X. Bismut and Lebeau [2] calculated the relation between the Quillen metric of η and ξ. With the help of this result, Gillet and Soulée [6] were able to proof a Riemann-Roch theorem in Arakelov geometry for the first Chern class of the direct image (see [12] for the theorem and some background information). This theorem was later proved by Faltings [3] for higher degrees.

The proof of the Riemann-Roch theorem uses a calculation of Gillet, Soulée and Zagier [5] of the torsion for the trivial line bundle on the complex projective spaces ℙ^{n}ℂ. This led Gillet and Soulée to conjecture this theorem, which was the initial motivation for [2]. This rather difficult calculation gives in particular the Gillet-Soulée R-genus, which appears explicitly in the theorem. This is the additive genus associated to the function

R(x) = ∑_{\begin{matrix}ℓ≥1 \\ odd \end{matrix}} (2ζ′(−ℓ) + ζ(−ℓ) ∑_{j=1}^{ℓ} \frac{1}{j})\frac{x^{ℓ}}{ℓ!} ,

where ζ is the Riemann zeta function. To obtain this series, one has to caculate the torsion of ℙ^{n}ℂ for every n.

Consider now a holomorphic isometry g of a hermitian vector bundle over a Kä ahler manifold. One can define an equivariant version of the torsion as an analogue to Ray’s [10] calculation of the real analytic torsion for lens spaces. Let M be a Kä ahler manifold of complex dimension n with holomorphic tangent bundle T M and Kä ahler metric g_{M}, ξ a hermitian vector bundle on M and \overline{∂} the Dolbeault operator acting on sections of Λ^{q}T^{∗(0,1)}M ⊗ ξ. We define the hermitian product on the vector space of smooth sections of Λ^{q}T^{∗(0,1)}M ⊗ ξ by

(η, η′) := ∫_{M}(η(x), η′(x))\frac{ω^{n}}{(2π)^{n}n!}

as in [5]. Consider the adjoint operator \overline{∂}^{∗} relative to this product and the Kodaira-Laplace operator

□_{q} := (\overline{∂} + \overline{∂}^{∗})^{2} : Γ(Λ^{q}T^{∗(0,1)}M ⊗ ξ) → Γ(Λ^{q}T^{∗(0,1)}M ⊗ ξ) .

Let g be a holomorphic isometry of M. Assume that the bundle and its hermitian metric are invariant under the induced action of g. Let Eig_{λ}(□_{q}) be the eigenspace of □_{q} corresponding to the eigenvalue λ and g^{∗} the of g induced action on Γ(Λ^{q}T^{∗(0,1)}M ⊗ ξ).

Consider the zeta function

Z(g, s) := ∑_{\begin{matrix}q>0 \\ λ∈ Spec⊓ ⊔_{q} \\ λ≠0 \end{matrix}}(−1)^{q+1}qλ^{−s} Trg_{| Eig_{λ}(□_{q})}^{∗}

for Re s ≫ 0. The equivariant torsion of M associated to the action of g is then defined as an exponential of minus half of the derivation at zero Z′(g, 0) of the holomorphic continuation of Z(g,.),

τ(g) := e^{− \frac{1}{2} Z′(g,0)} .

In this paper we present the calculation of the equivariant analytic torsion for all holomorphic bundles on ℙ^{1}ℂ and for the trivial line bundle on ℙ^{n}ℂ, where the projective spaces are equipped with the Fubini-Study metric. We consider only isometries with isolated fixed points. For a rotation by a angles ∈ π . ℚ, we obtain a closed expression involving the gamma function. For arbitrary angles it appears, as an infinite series, a part of the Gillet-Soulée R-function. This is relatively easy to calculate because the defining ζ -function Z has no singularities in contrast to the situation in [5]. The similarity between R and our result gives us the hope to an easier approach to their results.

For 0 < φ < 2π and Re s > 0, let ζ^{rot}(φ, s) be the Dirichlet series

ζ^{rot}(φ, s) := ∑_{k≥1} \frac{sin kφ}{k^{s}} .

Then ζ^{rot} can be seen as the imaginary part of a Lerch zeta function. We set R^{rot}(φ) := \frac{∂}{∂s}ζ^{rot}(φ, 0). One obtains easily by using classical results the

Proposition 1. R^{rot} is equal to

R^{rot}(φ) = \frac{C + ln φ}{φ} − ∑_{{ℓ≥1 \atop ℓ odd}} ζ′(−ℓ)(−1)^{\frac{ℓ+1}{2}} \frac{φ^{ℓ}}{ℓ!} .

If φ = 2π \frac{p}{q} with p, q ∈ ℕ, 0 < p < q, then

R^{rot}(φ) = −\frac{1}{2} ln q. cot \frac{φ}{2} + ∑_{ℓ=1}^{q−1} ln Γ (\frac{j}{q} ). sin jφ ,

where Γ is the gamma function and C = −Γ′(1) is Euler’s constant.

Let E := O(k_{1}) ⊕ … ⊕ O(k_{n}) be a holomorphic vector bundle on ℙ^{1}ℂ, equipped with the standard metric (i.e. the curvature of O(1) is the Fubini-Study Kä ahler form). By a theorem of Grothendieck, each holomorphic vector bundle on ℙ^{1}ℂ is of this form. Then we find

Theorem 2. The equivariant analytic torsion τ(E, φ) with respect to a rotation by an angle φ ∈]0, 2π[ of ℙ^{1}ℂ = S_{1/2}^{2} is given by

−2 ln τ(E, φ) = \frac{2R^{rot}(φ)}{sin \frac{φ}{2}}.∑_{j=1}^{n} cos(k_{j}+1)\frac{φ}{2}+∑_{j=1}^{n} ∑_{m=1}^{|k_{j}+1|} \frac{sin(2m − |k_{j} + 1|)\frac{φ}{2}}{sin \frac{φ}{2}} ln j .

We see in particular that the equivariant torsion τ gives already for the trivial line bundle O on ℙ^{1}ℂ the function

ln τ(O, φ) = cot \frac{φ}{2} . (i ∑_{\begin{matrix}ℓ≥1 \\ odd \end{matrix}} ζ′(−ℓ)\frac{(iφ)^{ℓ}}{ℓ!} − \frac{C + ln φ}{φ}) .

Let now Φ := ( ..._{0  iφ_{n+1}}^{iφ_{1}  0} ) be an element of the (canonical) maximal

Cartan subalgebras of 𝔰𝔲(n + 1), which defines an infinitesimal isometry on ℙ^{n}ℂ ≅ SU(n+1)/S(U(1)×U(n)). Assume that all the φ_{j} are distinct. Then we have

Theorem 3. The equivariant torsion τ(O, e^{Φ}) for the trivial line bundle O on ℙ^{n}ℂ is given by

−2 ln τ(O, e^{Φ}) = (−1)^{n} ∑_{{j,k=1 \atop j≠k}}^{n+1} 2iR^{rot}(φ_{j}−φ_{k}) ∏_{{ℓ=1 \atop ℓ≠k}}^{n+1}(e^{i(φ_{k}−φ_{ℓ})}−1)^{−1}− ln n! .

Proof (see [8]): The eigenvalues and eigenspaces for the Kodaira Laplacian for the trivial line bundle on ℙ^{n}ℂ were determined explicitly by Ikeda and Taniguchi [7]. If one regards ℙ^{n}ℂ as SU(n + 1)/S(U(1) × U(n)), the eigenspaces can be described by sums of irreducible representations of SU(n + 1). See for this idea also an article by Malliavin and Malliavin [9]. As Gillet and Soulée, we are using Ikeda’s and Taniguchi’s method and their results to determine explicitly the ζ -function Z in our case. The derivative at zero is then obtained by some calculations on Fourier series and determinants.

The function R^{rot} has a rather simple definition and hence a lot of special properties. We give here only a few of them.

Remark 4. The following identities hold

(1)  R^{rot}(φ) = −R^{rot}(2π − φ) ,
(2)  2R^{rot}(2φ) = R^{rot}(φ) + R^{rot}(π + φ) + ln 2 . cot φ ,
(3)  3R^{rot}(3φ) = R^{rot}(φ) + R^{rot}(\frac{2π}{3} + φ)

− R^{rot}(\frac{2π}{3} − φ) + \frac{3}{2} ln 3 . cot \frac{3φ}{2} ,

(4)  R^{rot}(π + φ) = ∫_{0}^{∞} ln x \frac{sinh φx}{sinh πx}dx .

Acknowledgement: The author would like to thank Prof. J.-M. Bismut for sharing the idea of this problem with him. This article is a part of the author’s thesis.

References

[1] J.-M. Bismut, H. Gillet, C. Soulée: Analytic torsion and holomorphic determinant bundles I, II, III, Comm. in Math. Physics 115 (1988), 49–78, 79–126, 301–351.

[2] J.-M. Bismut, G. Lebeau: Complex immersions and Quillen metrics, to appear in Publ. Math. IHES.

[3] G. Faltings: Lectures on the arithmetic Riemann-Roch theorem, Princeton University Press 1992.

[4] A Grothendieck: Sur la classification des fibrées holomorphes sur la sphèere de Riemann, Amer. J. Math. 79 (1956), 121–138.

[5] H. Gillet, C. Soulée: Analytic torsion and the arithmetic Todd genus, with an appendix by D. Zagier, Topology 30 (1991), 21–54.

[6] H. Gillet, C. Soulée: An arithmetic Riemann-Roch theorem, Preprint IHES/M/91/50.

[7] A. Ikeda, Y. Taniguchi: Spectra and Eigenforms of the Laplacian on S^{n} and ℙ^{n}ℂ, Osaka J. Math. 15 (1978), 515–546.

[8] K. Kö ohler: Equivariant analytic torsion on ℙ^{n}ℂ, to appear.

[9] M.-P. Malliavin, P. Malliavin: Diagonalisation du systèeme de de Rham-Hodge au-dessus d’un espace riemannien homogèene, Lect. Notes Math. 466 (1975), 135–146, Springer-Verlag.

[10] D.B. Ray: Reidemeister torsion and the Laplacian on lens spaces, Adv. in Math. 4 (1970), 109–126.

[11] D.B. Ray, I.M. Singer: Analytic torsion for complex manifolds, Ann. Math. 98 (1973), 154–177.

[12] C. Soulée, D. Abramovich, J.-F. Burnol, J. Kramer: Lectures on Arakelov Geometry, Cambridge studies in advanced math. 33, Cambridge University Press 1992.

Kai Kö ohler Déepartement de Mathéematique Universitée Paris-Sud F-91405 ORSAY